# neurips_2024.bib

### v4

@@ E.1 Architectures
In the following subsection, we describe the specific structure of the individual parts composing the purifier – in the three scenarios considered. As far as the classifier architectures are concerned, we redirect the reader to the original articles introducing those models (i.e.: [Cui2023Decoupled] for scenarios (a) and (b), [Wang2023Better] for scenario (c)).
@@
The original input is also fed to the classifier. The corresponding internal representation is extracted, preserving its layered structure. In order to improve the scalability of the method, only a subset of classifier layers is used instead of the whole internal representation. Specifically, for each block of the WideResNet architecture, only the first layers have been considered; two shortcut layers have also been added for good measure. The exact list of those layers is reported in Table 5.
@@
The sampler used for the generation of such latent variables $\bm{z}$, during the training of the purifier, is a reparameterised [Kingma2014Auto] Normal sampler $\bm{z}\sim\mathcal{N}(\mu,\sigma)$. During inference, $\bm{z}$ is sampled by reparameterisation from the i.i.d Standard Normal distribution $\bm{z}\sim\mathcal{N}(0,1)$ (i.e. from its original prior).
@@ E.2 Hyperparameters
Table 12: Hyperparameters for the attacks used for training and testing the purifier The Fgsm and Pdg attacks refer to the training phase (see subsection 4.4), whereas the Pgd+EoT attack [Lee2024Robust] refers to the robustness assessment pipeline. The entry CCE denotes the Categorical CrossEntropy loss function. The $\ell_{\infty}$ threat model is assumed, and all inputs are linearly rescaled within $[0.0,1.0]$ before the attack.
@@
All scenarios Sc. (a) Sc. (b) Sc. (c) Optimiser RAdam$+$Lookahead RAdam $\beta_{1}$ $0.9$ RAdam $\beta_{2}$ $0.999$ RAdam $\epsilon$ $10^{-8}$ RAdam Weight Decay None Lookahead averaging decay $0.8$ Lookahead steps $6$ Initial LR $5\times 10^{-9}$ Loss function CCE Sampled reconstructions per input $8$ Epochs $200$ $200$ $250$ LR warm-up epochs $25$ $25$ $31$ LR plateau epochs $25$ $25$ $31$ LR annealing epochs $150$ $250$ $188$ Plateau LR $0.064$ $0.064$ $0.0128$ Final LR $4.346\times 10^{-4}$ $4.346\times 10^{-4}$ $1.378\times 10^{-4}$ $\beta$ increase initial epoch $25$ $25$ $32$ $\beta$ increase final epoch $34$ $34$ $43$ Batch size $5120$ $2560$ $1024$ Adversarial batch fraction $0.5$ $0.15$ $0.01$
@@ Appendix F Additional tables
The best AT-based model for Cifar-10 is taken from [Cui2023Decoupled], whereas that for Cifar-100 from [Wang2023Better]. Both best purification-based models are taken from [Lin2024Robust].
@@
Scen. Dataset Best AT/Cl Best P/Cl (a) Cifar-10 0.9323 0.9082 (b) Cifar-100 0.7522 0.6973

### v5

@@ E.1 Architectures
In the following subsection, we describe the specific structure of the individual parts composing the purifier – in the three scenarios considered. As far as the classifier architectures are concerned, we redirect the reader to the original articles introducing those models (i.e., those by Cui et al. (2023) for scenarios (a) and (b), Wang et al. (2023) for scenario (c)).
@@
The original input is also fed to the classifier. The corresponding internal representation is extracted, preserving its layered structure. In order to improve the scalability of the method, only a subset of classifier layers is used instead of the whole internal representation. Specifically, for each block of the WideResNet architecture, only the first layers have been considered; two skip connections have also been added for good measure. The exact list of those layers is reported in Table 6.
@@
The sampler used for the generation of such latent variables $\bm{z}$, during the training of the purifier, is a reparameterised (Kingma & Welling, 2014) Normal sampler $\bm{z}\sim\mathcal{N}(\mu,\sigma)$. During inference, $\bm{z}$ is sampled by reparameterisation from the i.i.d Standard Normal distribution $\bm{z}\sim\mathcal{N}(0,1)$ (i.e. from its original prior).
@@ E.2 Hyperparameters
Table 13: Hyperparameters for the attacks used for training and testing the purifier The Fgsm and Pdg attacks refer to the training phase (see subsection 4.4), whereas the Pgd+EoT attack (Lee & Kim, 2024) refers to the robustness assessment pipeline. The entry CCE denotes the Categorical CrossEntropy loss function. The $\ell_{\infty}$ threat model is assumed, and all inputs are linearly rescaled within $[0.0,1.0]$ before the attack.
@@
All scenarios Sc. (a) Sc. (b) Sc. (c) Optimiser RAdam$+$Lookahead RAdam $\beta_{1}$ $0.9$ RAdam $\beta_{2}$ $0.999$ RAdam $\epsilon$ $10^{-8}$ RAdam Weight Decay None Lookahead averaging decay $0.8$ Lookahead steps $6$ Initial LR $5\times 10^{-9}$ Loss function CCE Sampled reconstructions per input $8$ Epochs $200$ $200$ $250$ LR warm-up epochs $25$ $25$ $31$ LR plateau epochs $25$ $25$ $31$ LR annealing epochs $150$ $150$ $188$ Plateau LR $0.064$ $0.064$ $0.0128$ Final LR $4.346\times 10^{-4}$ $4.346\times 10^{-4}$ $1.378\times 10^{-4}$ $\beta$ increase initial epoch $25$ $25$ $32$ $\beta$ increase final epoch $34$ $34$ $43$ Batch size $5120$ $2560$ $1024$ Adversarial batch fraction $0.5$ $0.15$ $0.01$
@@ Appendix F Additional tables
The best AT-based model for Cifar-10 is taken from Cui et al. (2023), whereas that for Cifar-100 from Wang et al. (2023). Both best purification-based models are taken from (Lin et al., 2024).
@@
Scen. Dataset Best AT/Cl Best P/Cl (a) Cifar-10 0.9368 0.9082 (b) Cifar-100 0.7522 0.6973